OpenStudy (anonymous):

Find the four arithmetic means between 3 and 28.

OpenStudy (anonymous):

nevermind i got the answer to be 8, 13, 18, 23

OpenStudy (igreen):

Yep, you got it.. We use the equation: 3 + 5d = 28 To find 'd', which is 5, then we add it to 8. 8 + 5 = 13 13 + 5 = 18 . . .